Partner A and Partner B have capital balances of $40,000 and $60,000, respectively, prior to the admission of Partner C. Partner C contributes $20,000 in exchange for a 20% interest in the partnership. The partnership agreement stipulates profits and losses are shared equally. What will be the balance in Partner A's capital account after the admission of Partner C?
Answer #1
Total equity=Total of old partner's capital+Contribution by new partner=40000+60000+20000=$ 120000
Share of C in Equity=Total equity*20%=120000*20%=$ 24000
Bonus received (paid) by new partner=Cash paid by new partner-Share in equity=20000-24000=$ -4000
Bonus to be paid to old partners equally
A 4000/2 2000
B 4000/2 2000
Balance in partner A's capital account=Beginning balance+Bonus received=40000+2000=$ 42000

> $42,000 is not correct.

Julia Hargrove Wed, Jan 12, 2022 7:18 PM